36-726 Statistical Practice Homework 03

1. Please read through the following materials, for discussion Monday Feb 19 in class:

Gopen, George D. and Swan, Judith A. (1990) The science of scientific writing. American Scientist, 78, 550–558. I don’t think this is in JSTOR but there is a copy in the “papers” area of

http://www.stat.cmu.edu/~brian/726/.
The “Language” slides at http://www.writing.eng.vt.edu/handbook/visuals.html.
The “Structure” slides at http://www.writing.eng.vt.edu/handbook/visuals.html.

2. Read the article, “hw03-example.pdf” that accompanies this hw on http://www.stat.cmu.edu/~brian/726/, and answer the following questions (bring your written answers to class, for discussion, Monday February 19):

(a) Write out all of the headings for sections, subsections, subsubsections, etc. What is good about this organization the paper? What could be improved? [write a sentence or two each]

(b) Can you find a paragraph whose sentences work really well in the sense of the Gopen et al. article above? List some things that make it easy to follow.

(c) Can you find a paragraph that really should be revised, according to Gopen et al.? Write a revision of it and tell what you tried to accomplish in your revision (correct any mechanics or other errors in your revision as well).